OpenStudy (republican31):

How many grams of water must be added to 3 grams of a 35% iodine solution in order to produce a 16% iodine solution? Round your answer to the nearest tenth. I had it set up like: 0.35(3)+0x=0.16(3-x) but I'm not sure if that's right.

OpenStudy (anonymous):

0.35*3=0.16*(x+3) = 3.5625 x= 3.6

OpenStudy (republican31):

Thank you!

OpenStudy (anonymous):

You're welcome! :)